OpenStudy (anonymous):

The x-intercept of 2x + 3y = 15 is (___, 0). A. 2 B. 3 C. 7.5 D. 5

jimthompson5910 (jim_thompson5910):

plug in y = 0, then solve for x 2x + 3y = 15 2x + 3*0 = 15 ... ... ... x = ???

OpenStudy (nurali):

2x + 3y = 15 2x + 3*0 = 15 2x=15 x=15/2 (x,y)=(15/2,0)